# Tracing Intricate Cues in Dialogue: Joint Graph Structure and Sentiment Dynamics for Multimodal Emotion Recognition

### v1

@@ Abstract
Multimodal emotion recognition in conversation (MERC) has garnered substantial research attention recently. Existing MERC methods face several challenges: (1) they fail to fully harness direct inter-modal cues, possibly leading to less-than-thorough cross-modal modeling; (2) they concurrently extract information from the same and different modalities at each network layer, potentially triggering conflicts from the fusion of multi-source data; (3) they lack the agility required to detect dynamic sentimental changes, perhaps resulting in inaccurate classification of utterances with abrupt sentiment shifts. To address these issues, a novel approach named GraphSmile is proposed for tracking intricate emotional cues in multimodal dialogues. GraphSmile comprises two key components, i.e., GSF and SDP modules. GSF ingeniously leverages graph structures to alternately assimilate inter-modal and intra-modal emotional dependencies layer by layer, adequately capturing cross-modal cues while effectively circumventing fusion conflicts. SDP is an auxiliary task to explicitly delineate the sentiment dynamics between utterances, promoting the model’s ability to distinguish sentimental discrepancies. Furthermore, GraphSmile is effortlessly applied to multimodal sentiment analysis in conversation (MSAC), forging a unified multimodal affective model capable of executing MERC and MSAC tasks. Empirical results on multiple benchmarks demonstrate that GraphSmile can handle complex emotional and sentimental patterns, significantly outperforming baseline models.
@@
publicationid: pubid: $00$ 2024 IEEE. Personal use is permitted, but republication/redistribution requires IEEE permission.
@@ I Introduction
In this work, we can effortlessly extend GraphSmile into a multitask model. Namely, GraphSmile is applied to the multimodal sentiment analysis in conversation (MSAC) task and achieves multitask joint optimization with the MERC task. Specifically, we initially incorporate a sentiment classifier, followed by the joint formulation of emotion classification loss, sentiment classification loss, and shift classification loss to construct a multitask learning objective, thereby synergistically optimizing network parameters. In this manner, GraphSmile is ultimately crafted into a unified multimodal affective model capable of executing MSAC and MERC tasks concurrently. In the domain of affective computing, emotions (e.g., happiness, surprise, neutrality, sadness, and anger) are typically regarded as more nuanced and fleeting psychological states, characterized by their situational and temporary nature, while sentiments (e.g., negative, neutral, positive) are understood as coarser and more enduring psychological inclinations, marked by their profundity and stability. While emotion and sentiment exhibit varying levels of granularity, duration, and characteristic, there exists an interrelation between them; happiness, for example, is typified as a positive sentiment, whereas sadness is reflective of a negative sentimental disposition. Thus, our co-optimization strategy not only facilitates knowledge transfer and complementary advantages between different granularity levels of affective tasks but also significantly enhances GraphSmile’s comprehensive perceptive capability for emotional and sentimental cues.
@@
  • A unified multitask multimodal affective model called GraphSmile is proposed for both MERC and MSAC tasks. GraphSmile integrates multiple losses to achieve multitask synergistic optimization while upgrading its ability to synthesize the perception of various cues.
@@ II-B Multimodal Sentiment Analysis
Multimodal sentiment analysis (MSA), due to its wide range of application prospects, has become a focal research topic in the field of affective computing. MSA is a multimodal task aimed at understanding and detecting human emotional tendencies by processing multimodal heterogeneous data, such as text, speech, and vision. The main challenge of the MSA task is how to effectively harness information from different modalities to complement each other, and a plethora of representative works [39, 40, 41, 42] have been proposed in succession. Existing mainstream works focus on enhancing the performance of MSA models by designing complex structures, interaction mechanisms, or fusion paradigms, which can be broadly categorized into tensor-based methods [43], attention-based methods [44, 41], and graph-based methods [45, 46]. These research trajectories of MSA tasks are distinctly different from those of the ERC tasks, as they do not take into account the conversational context, rendering them incompatible with ERC tasks. Although Hu et al. [11] have made effort to unify ERC and MSA, their approach to handling labels in the datasets is overly complex and cannot fully guarantee the correctness of universal labels. Divergent from the extant MSA methods, our GraphSmile is conceived as a multitask multimodal conversational affective model. In GraphSmile, we only need to merge emotional labels to effortlessly unify the ERC and MSA tasks, thereby facilitating an efficacious excavation of both emotional and sentimental cues in a concurrent manner.
@@ III-B1 Multimodal Dialogue Graphs
Edges: We establish relationships only between nodes of different modalities. In the graph $\mathcal{G}^{tv}$, a textual node $n_{i}^{t}$ is connected to all visual nodes; similarly, a visual node $n_{i}^{v}$ is connected to all textual nodes. To mitigate redundant connections [58, 18], we set up two sliding windows. Specifically, $n_{i}^{t}$ is associated with the past $P$ visual nodes (i.e., $n_{i-P}^{v},n_{i-P+1}^{v},\ldots,n_{i-1}^{v}$), and also with the future $F$ visual nodes (i.e., $n_{i+1}^{v},n_{i+2}^{v},\ldots,n_{i+F}^{v}$). The same connection way is applied to the visual node $n_{i}^{v}$. Additionally, different modal nodes (e.g., $n_{i}^{t}$ and $n_{i}^{v}$) from the same utterance are also connected.
@@
Following the construction manner of the graph $\mathcal{G}^{tv}$, a textual-acoustic dialogue graph $\mathcal{G}^{ta}$ and a visual-acoustic dialogue graph $\mathcal{G}^{va}$ can also be obtained. The majority of previous approaches [17, 18, 20], when creating cross-modal edges, only connect nodes from the same utterance, thereby neglecting the affective cues directly provided by other utterances. In contrast to these methods, in our construction of cross-modal edges, the current node is not only connected to nodes from the same utterance but also directly linked to nodes from different utterances. It is noteworthy that our multimodal dialogue graphs don’t establish edges between utterances in the same modality. Nonetheless, GraphSmile can still effortlessly capture intra-modal contextual affective cues through the unique graph convolution mechanism.
@@ III-B3 Example Explanation
$\begin{split}X^{(2)}&=AX^{(1)}\Theta^{(2)}\\ &=A(AX^{(0)}\Theta^{(1)})\Theta^{(2)}\\ &=A^{2}X^{(0)}\Theta^{(1)}\Theta^{(2)},\end{split}$ (5)
where $X^{(2)}$ and $\Theta^{(2)}$ represent the output and learnable parameter for the second layer, respectively. Despite the second layer’s features being predicated on those of the first layer, according to Eq. 5, the output of the second layer can also be directly computed from the initial input $X^{(0)}$. In our GSF module, since $A^{2}$ stores the second-hop neighborhood relationships of nodes, the significance of the second-layer graph convolution can be interpreted as nodes assimilating feature information from their second-hop neighbors. Referring to Fig. 2b, during the second-layer graph convolution, each node collects affective cues solely from nodes within the same modality. For instance, the textual node $n_{1}^{t}$ is connected only to nodes belonging to the textual modality, thereby collecting cues exclusively from intra-modal nodes (e.g., $n_{2}^{t}$, $n_{3}^{t}$); similarly, the visual node $n_{1}^{v}$ gathers affective cues only from visual nodes.
@@
$\begin{split}X^{(3)}&=AX^{(2)}\Theta^{(3)}\\ &=A(AX^{(1)}\Theta^{(2)})\Theta^{(3)}\\ &=A(A(AX^{(0)}\Theta^{(1)})\Theta^{(2)})\Theta^{(3)}\\ &=A^{3}X^{(0)}\Theta^{(1)}\Theta^{(2)}\Theta^{(3)},\end{split}$ (6)
where $X^{(3)}$ signifies the feature matrix resulting from the third layer and $\Theta^{(3)}$ represents the corresponding learnable parameter. Following Eq. 6, the third-layer graph convolution, akin to the second layer, can have its output directly computed from the initial features $X^{(0)}$. $A^{3}$ contains the third-hop neighborhood relationships, thus the implication of the third-layer is that nodes integrate feature information from their third-hop neighbors. As depicted in Fig. 2c, during the third-layer graph convolution, each node absorbs affective cues solely from nodes of another modality. For instance, similar to the first-layer convolution, the textual node $n_{1}^{t}$ is connected only to nodes of a different modality, i.e., visual nodes, and the visual node $n_{1}^{v}$ is connected only to textual nodes.
Through the aforementioned examples, it can be illustrated that our GSF module can alternately propagate inter-modal and intra-modal affective cues layer by layer, thereby preventing conflicts among multi-source information. Summarizing the patterns from the examples, to capture affective cues from more distant neighbors, we can extend to the $l$-th layer graph convolution:
@@ III-F Total Training Objective
where $\lambda_{s}$, $\lambda_{s}$, and $\lambda_{o}$ are trade-off parameters balancing these three losses; $\beta$ is the L2-regularizer weight, and $\Theta_{total}$ is the set of all learnable parameters.
@@ V-A Main Result for Emotion Recognition
The table also presents the F1 scores for each emotion. In the IEMOCAP (6-way) dataset, compared to other emotions, all models achieve the highest and second-highest scores on Sad and Excited, respectively, and the lowest on Happy. This indicates that Sad and Excited are relatively easier to distinguish than Happy. Relative to other models, our GraphSmile demonstrates a more balanced performance across all emotions, especially achieving the best score on the challenging emotion Happy. It can be seen that due to the reduction in the number of categories in the IEMOCAP dataset, these models have achieved satisfactory results on all emotions. Among these baselines, there is always one or more emotions with significantly lower scores than the others, indicating performance imbalance. In contrast, GraphSmile consistently achieves relatively high results with minimal variance across all emotions. Similar to the results on the IEMOCAP (6-way) dataset, our GraphSmile scores significantly higher on Happy than other models. In the MELD dataset, compared to other emotions, all models perform poorly on Fear and Disgust. The direct cause of this phenomenon is that these two emotions are extreme minority classes in the MELD dataset. Since samples labeled as Neutral constitute the majority of all samples, all models achieve the best F1 scores on Neutral. Compared to other models, SACL-LSTM and GraphSmile achieves the best and second-best scores on the minority class Disgust, respectively. In the CMU-MOSEI dataset, as each utterance may correspond to multiple emotions, we employ seven sentiment labels instead of emotion ones for our experiments. Table IV shows that the classification task for the CMU-MOSEI dataset is quite challenging, with all models struggling to distinguish each category, particularly failing to correctly identify Highly Negative. Compared to the baselines, GraphSmile achieves the best results on Negative, Weakly Positive, and Highly Positive.
@@ V-B Main Result for Sentiment Analysis
The proposed GraphSmile not only demonstrates exceptional performance in MERC but also successfully extends to the MSAC task. In Table V, we collect the experimental results of GraphSmile performing the MSAC task across multiple datasets. These results are obtained based on the same hyperparameter settings as in Table IV, showing that even without task-specific hyperparameter tuning, GraphSmile can still achieve commendable performance, especially surpassing all baselines on the IEMOCAP (4-way) dataset. The reduction in the number of categories leads to a significant performance improvement for all models on datasets other than IEMOCAP (4-way). After hyperparameter optimization specifically for the MSAC task, GraphSmile’s performance is further enhanced, outperforming all baselines on most datasets. Although SACL-LSTM achieves the best scores on the IEMOCAP (6-way) dataset, its performance on other datasets is not prominent. Similar to the results of the MERC task, DialogueCRN and SACL-LSTM show relatively weaker capabilities in capturing sentimental cues on the CMU-MOSEI dataset, with their performance significantly lower than other models. Additionally, the table lists the F1 score for each sentiment category in detail. In the IEMOCAP (6-way) dataset, the model scores lower in identifying Neutral, indicating that distinguishing Neutral is more challenging compared to other sentiments. A similar phenomenon is observed in the CMU-MOSEI dataset. In the IEMOCAP (4-way) dataset, the recognition of Negative is the most effective, while in the MELD dataset, the recognition score of Neutral is the highest. These results indicate that there are variations in the recognition difficulty for different datasets and sentiments.
@@ V-C Ablation Study
Multitask Learning: In ablation studies of multitask learning, we design three experiments, i.e., removing the sentiment classification task (No $\mathcal{L}_{s}$), removing the shift classification task (No $\mathcal{L}_{o}$), and removing both tasks simultaneously (No $\mathcal{L}_{s}$ $\mathcal{L}_{o}$). The results indicate that the removal of any single task adversely affects the overall performance of the model. In the MERC task, The performance decline attributable to the excision of $\mathcal{L}_{s}$ substantiates the capacity of the MSAC task to enhance the MERC task’s comprehension of emotions in a dialogue through knowledge complementation. Reciprocally, in the MSAC task, the performance decline induced by the ablation of $\mathcal{L}_{e}$ evidences that the MERC task can facilitate the MSAC task in understanding sentiments. In both tasks, the decrement in efficacy consequent to the elimination of $\mathcal{L}_{o}$ corroborates our SDP module can enable the model to capture the sentiment dynamics in a conversation, thus ensuring precise classification in scenarios with abrupt sentimental transitions. In addition, on the IEMOCAP dataset, the removal of $\mathcal{L}_{o}$ leads to a more significant performance drop compared to the removal of $\mathcal{L}_{s}$; conversely, on the MELD and CMU-MOSEI datasets, the observed phenomenon is the opposite of that on the IEMOCAP. This could be attributed to the abundance of long dialogues contained in the IEMOCAP dataset, whose sentimental changes is more complex relative to the other two datasets, thus necessitating a more urgent need for modeling sentiment dynamics.
@@
Modality Configuration: The empirical outcomes in modality ablation underscores a degradation in the model’s holistic performance following the removal of any modality, thereby revealing the indispensable complementarity of multimodal data in the MERC task. Specifically, across all examined datasets, the elimination of the textual modality (No t) engenders a markedly pronounced decrement in performance relative to other modalities. This revelation accentuates the dominant role of the textual modality in the MERC task, proffering more critical emotional information. Concurrently, it implies that the visual and acoustic modalities may carry relatively less emotional cues. Such disparities may stem from the noise embedded in visual and acoustic modalities, such as irrelevant background information, ambient sounds, and multiple facial expressions. Further analysis of the experimental results from the IEMOCAP dataset evinces that the exclusion of the visual modality (No v) exerts a relatively muted influence on the model’s efficacy compared to the excision of the acoustic modality (No a). This observation posits that in the IEMOCAP dataset, the visual modality contributes less to MERC than the other two modalities. Based on the results from the MELD and CMU-MOSEI datasets, it can be concluded that the contributions of the visual and acoustic modalities to the model’s performance on these two datasets are comparable, yet both are inferior to the textual modality.

### v2

@@ Abstract
Multimodal emotion recognition in conversation (MERC) has garnered substantial research attention recently. Existing MERC methods face several challenges: (1) they fail to fully harness direct inter-modal cues, possibly leading to less-than-thorough cross-modal modeling; (2) they concurrently extract information from the same and different modalities at each network layer, potentially triggering conflicts from the fusion of multi-source data; (3) they lack the agility required to detect dynamic sentimental changes, perhaps resulting in inaccurate classification of utterances with abrupt sentiment shifts. To address these issues, a novel approach named GraphSmile is proposed for tracking intricate emotional cues in multimodal dialogues. GraphSmile comprises two key components, i.e., GSF and SDP modules. GSF ingeniously leverages graph structures to alternately assimilate inter-modal and intra-modal emotional dependencies layer by layer, adequately capturing cross-modal cues while effectively circumventing fusion conflicts. SDP is an auxiliary task to explicitly delineate the sentiment dynamics between utterances, promoting the model’s ability to distinguish sentimental discrepancies. GraphSmile is effortlessly applied to multimodal sentiment analysis in conversation (MSAC), thus enabling simultaneous execution of MERC and MSAC tasks. Empirical results on multiple benchmarks demonstrate that GraphSmile can handle complex emotional and sentimental patterns, significantly outperforming baseline models.
@@
publicationid: pubid: 0162–8828  © 2025 IEEE. All rights reserved, including rights for text and data mining, and training of artificial intelligence and similar technologies. Personal use is permitted, but republication/redistribution requires IEEE permission. See https://www.ieee.org/publications/rights/index.html for more information.
@@ I Introduction
In this work, we can effortlessly extend GraphSmile into a multitask model. Namely, GraphSmile is applied to the multimodal sentiment analysis in conversation (MSAC) task and achieves multitask joint optimization with the MERC task. Specifically, we initially incorporate a sentiment classifier, followed by the joint formulation of emotion classification loss, sentiment classification loss, and shift classification loss to construct a multitask learning objective, thereby synergistically optimizing network parameters. In this manner, our GraphSmile is capable of efficiently executing both MSAC and MERC tasks. In the domain of affective computing, emotions (e.g., happiness, surprise, neutrality, sadness, and anger) are typically regarded as more nuanced and fleeting psychological states, characterized by their situational and temporary nature, while sentiments (e.g., negative, neutral, positive) are understood as coarser and more enduring psychological inclinations, marked by their profundity and stability. While emotion and sentiment exhibit varying levels of granularity, duration, and characteristic, there exists an interrelation between them; happiness, for example, is typified as a positive sentiment, whereas sadness is reflective of a negative sentimental disposition. Thus, our co-optimization strategy not only facilitates knowledge transfer and complementary advantages between different granularity levels of affective tasks but also significantly enhances GraphSmile’s comprehensive perceptive capability for emotional and sentimental cues.
@@
  • A multitask multimodal affective model that unites graph structure and sentiment dynamics, GraphSmile, is proposed for both MERC and MSAC tasks. GraphSmile exploits multiple losses to achieve multitask synergistic optimization while upgrading its ability to synthesize the perception of various cues.
@@ II-B Multimodal Sentiment Analysis
Multimodal sentiment analysis (MSA), due to its wide range of application prospects, has become a focal research topic in the field of affective computing. MSA is a multimodal task aimed at understanding and detecting human emotional tendencies by processing multimodal heterogeneous data, such as text, speech, and vision. The main challenge of the MSA task is how to effectively harness information from different modalities to complement each other, and a plethora of representative works [39, 40, 41, 42] have been proposed in succession. Existing mainstream works focus on enhancing the performance of MSA models by designing complex structures, interaction mechanisms, or fusion paradigms, which can be broadly categorized into tensor-based methods [39, 43], attention-based methods [44, 41], and graph-based methods [45, 46]. These research trajectories of MSA tasks are distinctly different from those of the ERC tasks, as they do not take into account the conversational context, rendering them incompatible with ERC tasks. Although Hu et al. [11] have made effort to unify ERC and MSA, their approach to handling labels in the datasets is overly complex and cannot fully guarantee the correctness of universal labels. Divergent from the extant MSA methods, our GraphSmile is conceived as a multitask multimodal conversational affective model. In GraphSmile, we only need to merge emotional labels to effortlessly unify the ERC and MSA tasks, thereby facilitating an efficacious excavation of both emotional and sentimental cues in a concurrent manner.
@@ III-B1 Multimodal Dialogue Graphs
Edges: We establish relationships only between nodes of different modalities. In the graph $\mathcal{G}^{tv}$, a textual node $n_{i}^{t}$ is connected to visual nodes; similarly, a visual node $n_{i}^{v}$ is connected to textual nodes. To mitigate redundant connections [58, 18], we set up two sliding windows. Specifically, $n_{i}^{t}$ is associated with the past $P$ visual nodes (i.e., $n_{i-P}^{v},n_{i-P+1}^{v},\ldots,n_{i-1}^{v}$), and also with the future $F$ visual nodes (i.e., $n_{i+1}^{v},n_{i+2}^{v},\ldots,n_{i+F}^{v}$). The same connection way is applied to the visual node $n_{i}^{v}$. Additionally, different modal nodes (e.g., $n_{i}^{t}$ and $n_{i}^{v}$) from the same utterance are also connected.
@@
Following the construction manner of the graph $\mathcal{G}^{tv}$, a textual-acoustic dialogue graph $\mathcal{G}^{ta}$ and a visual-acoustic dialogue graph $\mathcal{G}^{va}$ can also be obtained. The majority of previous approaches [17, 18, 20], when creating cross-modal edges, only connect nodes from the same utterance, thereby neglecting the affective cues directly provided by other utterances. In contrast to these methods, in our construction of cross-modal edges, the current node is not only connected to nodes from the same utterance but also directly linked to nodes from different utterances. It is noteworthy that our multimodal dialogue graphs don’t establish edges between utterances in the same modality. Nonetheless, GraphSmile can still effortlessly capture intra-modal contextual affective cues through the unique graph convolution mechanism. In Part III-B3, we explain how these intra-modal affective cues can be naturally extracted by graph convolution.
@@ III-B3 Example Explanation
$\begin{split}X^{(2)}&=AX^{(1)}\Theta^{(2)}\\ &=A(AX^{(0)}\Theta^{(1)})\Theta^{(2)}\\ &=A^{2}X^{(0)}\Theta^{(1)}\Theta^{(2)}.\end{split}$ (5)
Here, $X^{(2)}$ and $\Theta^{(2)}$ represent the output and learnable parameter for the second layer, respectively. Despite the second layer’s features being predicated on those of the first layer, according to Eq. 5, the output of the second layer can also be directly computed from the initial input $X^{(0)}$. In our GSF module, since $A^{2}$ stores the second-hop neighborhood relationships of nodes, the significance of the second-layer graph convolution can be interpreted as nodes assimilating feature information from their second-hop neighbors. Referring to Fig. 2b, during the second-layer graph convolution, each node collects affective cues solely from nodes within the same modality. For instance, the textual node $n_{1}^{t}$ is connected only to nodes belonging to the textual modality, thereby collecting cues exclusively from intra-modal nodes (e.g., $n_{2}^{t}$, $n_{3}^{t}$); similarly, the visual node $n_{1}^{v}$ gathers affective cues only from visual nodes. Thus, the above message passing process explains why our GraphSmile can still naturally capture intra-modal affective cues without creating intra-modal edges.
@@
$\begin{split}X^{(3)}&=AX^{(2)}\Theta^{(3)}\\ &=A(AX^{(1)}\Theta^{(2)})\Theta^{(3)}\\ &=A(A(AX^{(0)}\Theta^{(1)})\Theta^{(2)})\Theta^{(3)}\\ &=A^{3}X^{(0)}\Theta^{(1)}\Theta^{(2)}\Theta^{(3)}.\end{split}$ (6)
Here, $X^{(3)}$ signifies the feature matrix resulting from the third layer and $\Theta^{(3)}$ represents the corresponding learnable parameter. Following Eq. 6, the third-layer graph convolution, akin to the second layer, can have its output directly computed from the initial features $X^{(0)}$. $A^{3}$ contains the third-hop neighborhood relationships, thus the implication of the third-layer is that nodes integrate feature information from their third-hop neighbors. As depicted in Fig. 2c, during the third-layer graph convolution, each node absorbs affective cues solely from nodes of another modality. For instance, similar to the first-layer convolution, the textual node $n_{1}^{t}$ is connected only to nodes of a different modality, i.e., visual nodes, and the visual node $n_{1}^{v}$ is connected only to textual nodes.
Through the three examples mentioned above, it can be concluded that the first layer of graph convolution propagates affective cues across different modalities, the second layer of graph convolution conveys affective cues within the same modality, and the third layer of graph convolution once again captures inter-modal associative cues, and so on. Therefore, these examples illustrate that our GSF module can alternately propagate inter-modal and intra-modal affective cues layer by layer, thereby preventing conflicts among multi-source information. Summarizing the patterns from the examples, to capture affective cues from more distant neighbors, we can extend to the $l$-th layer graph convolution:
@@ III-F Total Training Objective
where $\lambda_{s}$ and $\lambda_{o}$ are trade-off parameters balancing these three losses; $\beta$ is the L2-regularizer weight, and $\Theta_{total}$ is the set of all learnable parameters.
@@ V-A Main Result for Emotion Recognition
The table also presents the F1 scores for each emotion. In the IEMOCAP (6-way) dataset, compared to other emotions, all models achieve the highest and second-highest scores on Sad and Excited, respectively, and the lowest on Happy. This indicates that Sad and Excited are relatively easier to distinguish than Happy. Relative to other models, our GraphSmile demonstrates a more balanced performance across all emotions, especially achieving the best score on the challenging emotion Happy. As can be seen from the results of the IEMOCAP (4-way) dataset, due to the reduction in the number of categories in the dataset, these models have achieved satisfactory results on all emotions. Among these baselines (with the exception of MMGCN), there is always one or more emotions with significantly lower scores than the others, indicating performance imbalance. In contrast, GraphSmile consistently achieves relatively high results across all emotions. Similar to the results on the IEMOCAP (6-way) dataset, our GraphSmile scores significantly higher on Happy than other models. In the MELD dataset, compared to other emotions, all models perform poorly on Fear and Disgust. The direct cause of this phenomenon is that these two emotions are extreme minority classes in the MELD dataset. Since samples labeled as Neutral constitute the majority of all samples, all models achieve the best F1 scores on Neutral. Compared to other models, SACL-LSTM and GraphSmile achieves the best and second-best scores on the minority class Disgust, respectively. In the CMU-MOSEI dataset, as each utterance may correspond to multiple emotions, we employ seven sentiment labels instead of emotion ones for our experiments. Table IV shows that the classification task for the CMU-MOSEI dataset is quite challenging, with all models struggling to distinguish each category, particularly failing to correctly identify Highly Negative. Compared to the baselines, GraphSmile achieves the best results on Negative, Weakly Positive, and Highly Positive.
@@ V-B Main Result for Sentiment Analysis
The proposed GraphSmile not only demonstrates exceptional performance in MERC but also successfully extends to the MSAC task. In Table V, we collect the experimental results of GraphSmile performing the MSAC task across multiple datasets. The results obtained based on the hyperparameter settings in Table IV show that even without task-specific hyperparameter tuning, GraphSmile can still achieve commendable performance, especially surpassing all baselines on the IEMOCAP (4-way) dataset. The reduction in the number of categories leads to a significant performance improvement for all models on datasets other than IEMOCAP (4-way). After hyperparameter optimization specifically for the MSAC task, GraphSmile’s performance is further enhanced, outperforming all baselines on most datasets. Although SACL-LSTM achieves the best scores on the IEMOCAP (6-way) dataset, its performance on other datasets is not prominent. Similar to the results of the MERC task, DialogueCRN and SACL-LSTM show relatively weaker capabilities in capturing sentimental cues on the CMU-MOSEI dataset, with their performance significantly lower than other models. Additionally, the table lists the F1 score for each sentiment category in detail. In the IEMOCAP (6-way) dataset, the model scores lower in identifying Neutral, indicating that distinguishing Neutral is more challenging compared to other sentiments. A similar phenomenon is observed in the CMU-MOSEI dataset. In the IEMOCAP (4-way) dataset, the recognition of Negative is the most effective, while in the MELD dataset, the recognition score of Neutral is the highest. These results indicate that there are variations in the recognition difficulty for different datasets and sentiments.
@@ V-C Ablation Study
Multitask Learning: In ablation studies of multitask learning, we design three experiments, i.e., removing the sentiment classification task (No $\mathcal{L}_{s}$), removing the shift classification task (No $\mathcal{L}_{o}$), and removing both tasks simultaneously (No $\mathcal{L}_{s}$ $\mathcal{L}_{o}$). The results indicate that the removal of any single task adversely affects the overall performance of the model. In the MERC task, the performance degradation attributable to the excision of $\mathcal{L}_{s}$ substantiates the capacity of the MSAC task to enhance the MERC task’s comprehension of emotions in a dialogue through knowledge complementation. Reciprocally, in the MSAC task, the performance decline induced by the ablation of $\mathcal{L}_{e}$ evidences that the MERC task can facilitate the MSAC task in understanding sentiments. In both tasks, the decrement in efficacy consequent to the elimination of $\mathcal{L}_{o}$ corroborates our SDP module can enable the model to capture the sentiment dynamics in a conversation, thus ensuring precise classification in scenarios with abrupt sentimental transitions. In addition, on the IEMOCAP dataset, the removal of $\mathcal{L}_{o}$ leads to a more significant performance drop compared to the removal of $\mathcal{L}_{s}$; conversely, on the MELD and CMU-MOSEI datasets, the observed phenomenon is the opposite of that on the IEMOCAP. This could be attributed to the abundance of long dialogues contained in the IEMOCAP dataset, whose sentimental changes is more complex relative to the other two datasets, thus necessitating a more urgent need for modeling sentiment dynamics.
@@
Modality Configuration: The empirical outcomes in modality ablation underscores a degradation in the model’s holistic performance following the removal of any modality, thereby revealing the indispensable complementarity of multimodal data in the MERC task. Specifically, across all examined datasets, the elimination of the textual modality (No Textual) engenders a markedly pronounced decrement in performance relative to other modalities. This revelation accentuates the dominant role of the textual modality in the MERC task, proffering more critical emotional information. Concurrently, it implies that the visual and acoustic modalities may carry relatively less emotional cues. Such disparities may stem from the noise embedded in visual and acoustic modalities, such as irrelevant background information, ambient sounds, and multiple facial expressions. Further analysis of the experimental results from the IEMOCAP dataset evinces that the exclusion of the visual modality (No Visual) exerts a relatively muted influence on the model’s efficacy compared to the excision of the acoustic modality (No Acoustic). This observation posits that in the IEMOCAP dataset, the visual modality contributes less to MERC than the other two modalities. Based on the results from the MELD and CMU-MOSEI datasets, it can be concluded that the contributions of the visual and acoustic modalities to the model’s performance on these two datasets are comparable, yet both are inferior to the textual modality.